Determine the mass of a moving object with a velocity of 35 m/s and a kinetic energy of 58 J.


1
Expert's answer
2022-02-07T14:02:51-0500

Given:

"v=35\:\rm m/s"

"E_k=58\:\rm J"


The kinetic energy is given by

"E_k=\frac{mv^2}{2}"

The mass of object

"m=\frac{2E_k}{v^2}=\frac{2*58\:\rm J}{(35\:\rm m/s)^2}=0.095\:\rm kg"